Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

flip

Expert replies
by ruplun » Fri Jul 29, 2011 5:32 am
Kevin flips a coin 4 times .What is the probability that he gets heads on at least one of the 4 flips?
Join the discussion
Source: — Problem Solving |

by Anurag@Gurome » Fri Jul 29, 2011 5:40 am
ruplun wrote:Kevin flips a coin 4 times .What is the probability that he gets heads on at least one of the 4 flips?
Probability that Kevin gets heads on none of the 4 flips = (1/2)^4 = 1/16
Therefore, probability that Kevin gets heads on at least one of the 4 flips = 1 - 1/16 = [spoiler]15/16[/spoiler]
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by Brent@GMATPrepNow » Fri Jul 29, 2011 5:41 am
ruplun wrote:Kevin flips a coin 4 times .What is the probability that he gets heads on at least one of the 4 flips?
We can solve this, we'll using the complement (complements are often useful for questions involving "at least").

So, P(at least 1 head) = 1 - P(no heads)
P(no heads) = P(all tails)

Aside: we'll use T1 to denote tails on first flip, T2 for tails on 2nd flip and so on.

P(all tails) = P(T1 AND T2 AND T3 AND T4)
P(all tails) = P(T1) x P(T2) x P(T3) x P(T4)
P(all tails) = 1/2 x 1/2 x 1/2 x 1/2
P(all tails) = 1/16


So, P(at least 1 head) = 1 - P(no heads)
P(at least 1 head) = 1 - 1/16
P(at least 1 head) = 15/16

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion